# Accelerating Equity: Overcoming the Gender Gap in VC Funding

### v1

@@ 3.3 The Gender Gap in Startup Performance
Table 4 summarizes startup performance post-accelerator. Women-founded startups had a similar likelihood of securing VC funding within one year of graduation compared to men-founded startups (0.429 vs. 0.443) but raised smaller amounts on average (\$0.352 m vs. \$0.633 m). In terms of performance, women-founded startups had lower failure rates within one year but higher failure rates in the long run. [^9] The patterns in Figure 1 hold also with different definition of women-funded startup, as we show in Appendix A.2.
Table 5 breaks down fundraising performance by VC funding amount. While women-founded startups performed similarly at lower funding levels, significantly fewer raised more than \$2 million.
A bootstrapping test (Appendix Table A2) shows that among funded startups, the difference in mean log-investment size between men- and women-founded startups is not significantly greater than zero (p-values: 0.293 within one year, 0.160 within five years). However, the log-investment size distribution for men-founded startups exhibits significantly higher variance and right skewness. This suggests that male founders are more likely to secure exceptionally large funding amounts, whereas female founders face greater challenges in obtaining substantial investments.
@@
All Women Founded Men Founded Mean St. dev. Mean St. dev. Mean St. dev. VC Investments Fraction of VC funded startups (1 yr since grad) 0.442 0.497 0.429 0.499 0.443 0.497 Fraction of VC funded startups (5 yr since grad) 0.510 0.500 0.429 0.499 0.517 0.500 Investment size (m\$) (1 yr since grad) 0.609 1.442 0.352 0.540 0.633 1.497 Investment size (m\$) (5 yr since grad 6.544 34.406 1.341 2.769 7.031 35.935 Operation Status Fractions of startups that failed (1 yr since grad) 0.045 0.207 0.032 0.177 0.046 0.210 Fractions of startups that failed (5 yr since grad) 0.345 0.476 0.444 0.501 0.336 0.473 Fractions of startups that were acquired (5 yr since grad) 0.228 0.420 0.206 0.408 0.230 0.421 Observations 736 63 673
  • Notes: This table presents the performance of startups after graduating from an accelerator. By investment size, we denote the average amount of capital received from VCs.
@@
Funded Startups Funded Amount Number of Yes No >1m >2m >5m Startups One year after graduation Men founded 375 298 119 46 22 673 (%) (55.7) (44.3) (17.7) (6.8) (3.3) (100.0) Women founded 36 27 9 1 0 63 (%) (57.1) (42.9) (14.3) (1.6) (0.0) (100.0) All 411 325 128 47 22 736 (%) (55.8) (44.2) (17.4) (6.4) (3.0) (100.0) Five years after graduation Men founded 325 348 223 174 123 673 (%) (48.3) (51.7) (33.1) (25.9) (18.3) (100.0) Women founded 36 27 18 12 5 63 (%) (57.1) (42.9) (28.6) (19.0) (7.9) (100.0) All 361 375 241 186 128 736 (%) (49.0) (51.0) (32.7) (25.3) (17.4) (100.0)
  • Notes: This table shows the number and percentage of startups that received VC funding: not funded, funded, funded with more than 1 million US\$, funded with more than 2 million US\$, and funded with more than 5 million US\$.
@@ 4 A Matching Model for VCs and Startups
We build a model to discern accelerator and startup quality and to measure the extent of discrimination in this industry (Sørensen,, 2007). The model features assortative matching, so that high-quality startups and accelerators are more likely to match during admission. We also allow other types of complementarties (e.g., geographic proximity) to affect a match probability. [^10] Hallen et al., (2020) provide evidence for assortative matching between startups and accelerators. Thus, part of a startup’s unobserved quality can be inferred from the quality of the accelerator it attends.
@@ 4.1 A Matching Model with Non-transferable Utility
We model accelerator admissions as a two-sided matching game with non-transferable utility, following Roth and Sotomayor, (1990). Each accelerator-startup match generates a joint match value, which is split according to a fixed equity share. This value depends on observed and latent characteristics of both accelerators and startups, including quality measures and complementarities. The equity share is exogenous and identical across matches. This assumption is supported by the data: the average accelerator equity share is $6.2\%$ with a standard deviation of $1\%$, and all matches within an accelerator have the same equity share.
@@
Market definition. Our market definition is quite flexible as each market comprises all U.S. accelerators and their program participants within a six-month period, beginning in January 2008 and ending in June 2011. This definition imposes fewer geographical constraints than Sørensen, (2007)—who uses smaller regional areas—given the prevalence of out-of-state matches.
We adopt a semiannual market frequency to align with accelerators that run two programs per year, such as Y Combinator. This avoids self-overlapping while ensuring market independence. Sensitivity checks confirm that adjusting market windows does not alter qualitative results. We treat consecutive markets as independent, following Sørensen, (2007) and Fox, (2018), and omit dynamic considerations, as startups typically attend accelerators only once.
The information structure. We assume that all startups seeking accelerator programs are aware of all available programs within the same market. This is a standard assumption in the literature, justified by the widespread availability of public announcements online months before admissions begin. However, we do not assume that accelerators have complete knowledge of all potential startups in the market.
Complementarities in portfolio selection. To ensure the existence of a stable match, we abstract from complementarities in accelerator portfolio selection, assuming that an accelerator’s preference for one startup is independent of its preference for others. [^11] The evidence in Section 3 suggests that accelerators’ and VCs’ evaluations of startups are closely aligned, thus supporting our assumption that both VCs and accelerators view startup characteristics, those that are unobserved by researchers, in a similar way. Therefore, we can infer accelerator preferences via admission outcomes to approximate VC assessments of these unobserved characteristics. This assumption is standard in empirical two-sided matching models (e.g., Mindruta et al.,, 2016; Akkus et al.,, 2020; Pan,, 2017; Honoré and Ganco,, 2020; Akkus et al.,, 2016; Fox,, 2018). In practice, accelerators face competition, making deliberate portfolio construction difficult, and it is not uncommon for direct competitors to be admitted to the same cohort (Stross,, 2012).
@@ 4.1.1 Setup of the Theoretical Model
First stage (matching). Let $A$ be the set of accelerators and $S$ the set of startups in a market. A potential match is denoted by $(a,s)$ for $a\in A$ and $s\in S$. For the pair $(a,s)$, $a$ and $s$ share a total match value $U_{as}$. Let $U_{as}^{a}$ and $U_{as}^{s}$ be the payoffs for $a$ and $s$ from match $(a,s)$, respectively. We have:
@@
A matching is a function $\mu$ from the set of startups $S$ to the set of accelerators $A$. The equality $\mu(s)=a$ indicates that $s$ is matched to $a$ under matching $\mu$. The solution concept relies on the “no-blocking condition:” A pair $(a,s)$ is blocking for $\mu$ if its two entries are not matched but prefer each other over one of their current match(es). Mathematically, if $(a,s)$ is a blocking pair for $\mu$, then we have $\mu(s)\neq a$ and simultaneously,
$\begin{cases}U_{as}>U_{\mu(s),s}\\ U_{as}>\min_{s^{\prime}\in\mu^{-1}(a)}U_{as^{\prime}}\end{cases}.$
That is, $s$ prefers $a$ to its current match $\mu$, and $a$ prefers $s$ to at least one of its current matches in $\mu^{-1}(a)$. A matching $\mu$ is “stable’ if there is no blocking pair. Proposition 1 in Sørensen, (2007) applies to our framework and ensures the existence of a unique stable matching $\mu$ [^12] Without a unique equilibrium, the statistical likelihood function lacks a well-defined structure, and without a straightforward equilibrium condition, the empirical model becomes intractable (Bresnahan and Reiss,, 1991).
Let the observed covariates of $a$ and $s$ be $X^{as}$. Given the distribution of $\epsilon^{as}$, our first-stage estimator recovers the matching parameters $\beta$ in the expression
$U_{as}=X^{as}\beta+\epsilon^{as}.$ (1)
The term $\epsilon^{as}$ contains idiosyncratic unobserved factors that affect the match value for the pair $(a,s)$.
Second stage (longer-term performance). In the second stage, post-matching performance, $Y^{as}$, is modeled as
$Y^{as}=X^{as}\alpha+\eta^{as}.$ (2)
Here, $Y$ can be any startup’s performance after it finishes the accelerator program. In particular, the coefficient of gender is our parameter of interest, and it lies in $\alpha$.
The second-stage error term is $\eta^{as}$ correlated with $\epsilon^{as}$. This is because the VC’s decision $Y$ is also correlated with the unobserved match quality $\epsilon^{as}$, which is observed by the accelerator managers and the startups during the admission process. For each potential pair $(a,s)$, we model the random vector $(\epsilon^{as},\eta^{as})$ using a bivariate normal distribution. Without loss of generality in our context, we normalize the variances so that
$\begin{bmatrix}\epsilon^{as}\\ \eta^{as}\end{bmatrix}\sim\mathcal{N}\left(0,\begin{bmatrix}1&\rho\sigma\\ \rho\sigma&\sigma^{2}\end{bmatrix}\right).$
@@ 4.2.1 First Stage: Estimating the Matching Model
We express the likelihood function using potential blocking pairs. For convenience, we denote by $\bar{U}_{as}:=X^{as}\beta$ the deterministic component of the value from the potential pair $(a,s)$. For any pair $(a,s^{\prime})$ where $\mu(s^{\prime})\neq a$, it is not a blocking pair of $\mu$ if
@@
do not hold simultaneously. Define
@@
Note that $\underline{U}_{as^{\prime}}$ depends on $\epsilon^{as}$ for $s\in\mu^{-1}(a)$ and $\epsilon^{a^{\prime}s^{\prime}}$ for $a^{\prime}=\mu(s^{\prime})$. Therefore, given the unobservables $\epsilon^{as}$ for each observed pair $(a,s)$ that satisfies $a=\mu(s)$, the probability that $\mu$ is the equilibrium is
@@
where $\Phi$ is the c.d.f. of a standard normal distribution. Since this product readily integrates out all unmatched pairs $(a,s^{\prime})$, the overall likelihood of an observed matching $\mu$ is therefore
@@
By considering $\epsilon^{as}$ as latent for all matched pairs $(a,s)$ where $a=\mu(s)$, we can obtain the maximum likelihood estimates for the parameters $\beta$ using a simulated likelihood approach. The confidence intervals for the parameters are obtained through bootstrapping (500 repetitions) because, in this matching framework, it is theoretically unclear whether the number of accelerators or the number of startups must go to infinity for the standard asymptotic result to hold.
@@ 4.2.2 Second Stage: Estimating the Performance of Startup Graduates
In the second-stage analysis, we study ex post outcomes $Y^{as}$ through (2). However, $\eta^{as}$ is not independent of $X^{as}$ due to its correlation with $\epsilon^{as}$. For example, when $a$ and $s$ are distant from each other, they can still form a match if the realized $\epsilon^{as}$ is large enough. Therefore, among the realized matches, distance correlates with $\epsilon^{as}$. As a result, since $\epsilon^{as}$ correlates with $\eta^{as}$, $\eta^{as}$ correlates with $X^{as}$.
To have consistent results, we must control for $\mathbb{E}[\eta^{as}|\mu,X]$, the conditional expectation of each $\eta^{as}$ given the realized matching with all observable characteristics $X$ in the market. In other words, suppose that we observed $\widehat{\mathbb{E}[\eta^{as}|\mu,X]}$, we could use it in (2) to correct for endogeneity problems through a control function approach:
$Y^{as}=X^{as}\alpha+\widehat{\mathbb{E}[\eta^{as}|\mu,X]}+\delta$ (3)
where $\delta:=\eta^{as}-\widehat{\mathbb{E}[\eta^{as}|\mu,X]}$. Since the error term $\delta$ now has expectation zero given $X^{as}$ (i.e., $X^{as}$ is controlled for in $X$), this regression can be estimated with no bias through the ordinary least squares (OLS) method, according to:
Proposition 1
$\mathbb{E}[\eta^{as}|\mu,X]$ is a scalar multiple of $\mathbb{E}[\epsilon^{as}|\mu,X]$, i.e., they are collinear.
Proof. See Appendix section A.4. $\Box$
The proposition finds that we can use the observed $\mathbb{E}[\epsilon^{as}|\mu,X]$ in place of the unobserved $\mathbb{E}[\epsilon^{as}|\mu,X]$ in (3) to obtain an unbiased least squares estimate for $\alpha$. We estimate $\mathbb{E}[\epsilon^{as}|\mu,X]$ by taking the average of the simulated conditional distribution of $\mathbb{E}[\epsilon^{as}|\mu,X]$ from the matching model. The standard errors for $\alpha$ is obtained through bootstrapping. [^13] In the case where the outcome of interest is a binary variable, we simplify the computation using the linear probability model in the regression. We check whether the estimates from a probit model are qualitatively similar.
@@ 4.3 Observables
Gender gap. To examine the gender gap, we define Female Founder as an indicator for whether at least one startup founder is female. While analyzing all-women founding teams is of interest, the limited number of such cases constrains our analysis. This measure aligns with the divergence trend observed in Pitchbook data, where gender disparities are likely more pronounced for all-women teams (Box and Segerlind,, 2018).
Team member characteristics. We collect other characteristics of the founding members with information collected from Linkedin and their personal websites. To capture prior entrepreneurship experience, we define No Serial Founder as an indicator equal to one if no founder has prior startup experience. General work experience, proxied by age, serves as a quality signal, which we examine through Average Age of Founding Team. We also include indicators for educational background: At least One Graduate Degree, At least One PhD Degree, and At least One Engr/Sci Degree. Additionally, Founding Team Size captures the number of founders. Industry differences are accounted for using six-digit 2012 NAICS codes.
Startup characteristics. Longer-established startups differ from newly founded ones, as they are more likely to have a defined business model, customer base, and experienced founders, all of which facilitate investor attraction. To account for this, we include Startup Age, defined as the number of years since founding prior to accelerator admission.
Accelerator characteristics. We capture accelerator quality variation with log(Cohort Size), while Accelerator Experience measures the years an accelerator has operated before the current cohort. Accelerator in Startup Hubs indicates whether the accelerator is in MA, CA, or NY. [^14] Data limitations prevent controlling for detailed location fixed effects. To account for potential gender-related differences, ‘All-Men’ Accelerator equals 1 if the accelerator has no female founder.
Second-stage performance variables. We measure startup performance using total VC investment received within one year and five years after demo day. Following the literature, we also track survival status (failed or still operating) and acquisition status (acquired or not) by the fifth year (Hockberg et al.,, 2007; Gompers et al.,, 2010; Ewens and Rhodes-Kropf,, 2015). Since very few accelerator graduates have completed an IPO, we do not consider IPOs as a measure of performance.

### v2

@@ 3.3 The Gender Gap in Startup Performance
We define a startup as “women-founded” if at least one of its founders is female. While this definition is broad, it is necessary given the scarcity of startups with all-women founding teams. Using data on the universe of U.S. startups from Pitchbook, (2021, 2022), Appendix A.2 shows that similar patterns to the ones we discuss in this paper emerge under alternative definitions of “women-founded,” including those restricted to all-women teams. As a result, our estimates should be interpreted as a lower bound on the true gender gap in startup founding—both due to the inclusive definition of “women-founded” and because accelerators tend to attract the most promising startups.
Table 3 compares the performance post-accelerator of women- and men-founded startups. It finds that women-founded startups had a similar likelihood of securing VC funding within one year of graduation compared to men-founded startups (0.429 vs. 0.443) but raised smaller amounts on average (\$0.352 m vs. \$0.633 m). In terms of performance, women-founded startups had lower failure rates within one year but higher failure rates in the long run.
Table 4 further breaks down fundraising performance by VC funding amount. Note that we observe no difference in the fraction of funded startups across genders (Columns 1 and 2). However, while women-founded startups performed similarly to men-founded startups in percentage at lower funding levels (Column 3), significantly fewer of them raised more than \$2 million both one year and five years from graduation (Columns 4 and 5). [^8] To confirm this, a bootstrapping test in Appendix Table A2 shows no significant difference in mean log-investment size between men- and women-founded startups (p-values: 0.293 at one year, 0.160 at five years). However, the distribution for men-founded startups has significantly higher variance and right skewness, indicating greater access to exceptionally large funding.
@@
All Women Founded Men Founded VC Investments (1 and 5 years) Fraction VC funded (1 year) 0.442 0.429 0.443 Fraction VC funded (5 years) 0.510 0.429 0.517 Investment size (million \$, 1 year) 0.609 0.352 0.633 Investment size (million \$, 5 years) 6.544 1.341 7.031 Operational Status (1 and 5 years) Failure rate (1 year) 0.045 0.032 0.046 Failure rate (5 years) 0.345 0.444 0.336 Acquisition rate (5 years) 0.228 0.206 0.230 Observations 736 63 673 Notes: This table reports average startup outcomes after graduating from an accelerator. Investment size reflects average VC capital raised.
@@
(1) (2) (3) (4) (5) (6) Funded Startups Funded Amount Number of Yes No ¿1m ¿2m ¿5m Startups One year after graduation Men founded 375 298 119 46 22 673 (%) (55.7) (44.3) (17.7) (6.8) (3.3) (100.0) Women founded 36 27 9 1 0 63 (%) (57.1) (42.9) (14.3) (1.6) (0.0) (100.0) All 411 325 128 47 22 736 (%) (55.8) (44.2) (17.4) (6.4) (3.0) (100.0) Five years after graduation Men founded 325 348 223 174 123 673 (%) (48.3) (51.7) (33.1) (25.9) (18.3) (100.0) Women founded 36 27 18 12 5 63 (%) (57.1) (42.9) (28.6) (19.0) (7.9) (100.0) All 361 375 241 186 128 736 (%) (49.0) (51.0) (32.7) (25.3) (17.4) (100.0) Notes: This table shows the number and percentage of startups that received VC funding: not funded, funded, funded with more than 1 million US\$, funded with more than 2 million US\$, and funded with more than 5 million US\$.
@@ 4 A Matching Model for VCs and Startups
We build a model to discern accelerator and startup quality and to measure the extent of discrimination in this industry. The model features assortative matching, so that high-quality startups and accelerators are more likely to match during admission. We also allow other types of complementarties (e.g., geographic proximity) to affect a match probability. Thus, part of a startup’s unobserved quality can be inferred from the quality of the accelerator it attends.
@@ 4.1 A Matching Model with Non-transferable Utility
[…]
@@
Market definition.
We define each market as the set of all U.S. accelerators and their program participants within a six-month cohort, beginning in January 2008 and ending in June 2011. The first market includes all matches formed between January and June 2008, the second spans July to December 2008, and so on, through successive six-month periods. This temporal definition avoids imposing strong geographical constraints, unlike prior work—for example, Sørensen, (2007) segments the entire U.S. market into smaller local regions. Given the high incidence of out-of-state matches between startups and accelerators, our approach provides a more flexible and empirically relevant market structure.
We adopt a semiannual market frequency to align with accelerators that run two programs per year, such as Y Combinator. We treat consecutive markets as independent, following Sørensen, (2007) and Fox, (2018), and omit dynamic considerations, as startups typically attend accelerators only once. Sensitivity checks confirm that adjusting market windows does not alter the results qualitatively.
The information structure.
We assume that all startups seeking accelerator programs are aware of all available programs within the same market. This is a standard assumption in the literature, justified by the widespread availability of public announcements online months before admissions begin. However, we do not assume that accelerators have complete knowledge of all potential startups in the market.
Complementarities in portfolio selection.
To ensure the existence of a stable match, we abstract from complementarities in accelerator portfolio selection, assuming that an accelerator’s preference for one startup is independent of its preference for others. Also this assumption is standard in empirical two-sided matching models (e.g., Mindruta et al., , 2016; Akkus et al., , 2020; Pan, , 2017; Honoré and Ganco, , 2020; Akkus et al., , 2016; Fox, , 2018). In practice, accelerators face competition, making deliberate portfolio construction difficult, and it is not uncommon for direct competitors to be admitted to the same cohort (Stross, , 2012).
@@ 4.1.1 Setup of the Theoretical Model
The novelty of our approach lies in keeping the two estimation steps separate. While joint estimation is statistically more efficient, it becomes computationally infeasible in large markets, as the joint likelihood lacks a closed-form expression. As a result, researchers must simulate the distribution over all possible matches, typically restricting analysis to small, local markets (Sørensen, , 2007). However, this limitation may preclude reallocation across distant markets—a key mechanism in our analysis of the gender gap and a quantitatively important channel, with 27% of startups relocating to a different state. By contrast, our two-step approach yields an analytical expression for the matching likelihood in the first step, enabling estimation in settings with many potential matches. In the second step, it allows us to account for match quality using a simple control function approach to study startup performance.
First stage (matching).
[…]
@@
[…]
$\begin{cases}U_{as}>U_{\mu(s),s}\\ U_{as}>\min_{s^{\prime}\in\mu^{-1}(a)}U_{as^{\prime}}.\end{cases}$
That is, $s$ prefers $a$ to its current match $\mu$, and $a$ prefers $s$ to at least one of its current matches in $\mu^{-1}(a)$. A matching $\mu$ is “stable’ if there is no blocking pair. In this context, we can demonstrate that:
Proposition 1.
The accelerator-startup matching model defined under the preferences in 4.1.1 admits a unique equilibrium.
The proof of Proposition 1 follows from applying Proposition 1 in Sørensen, (2007) to our framework. Establishing the existence of a unique stable matching $\mu$ is crucial: without it, the statistical likelihood function is ill-defined, and the absence of a clear equilibrium condition renders the empirical model intractable (Bresnahan and Reiss, , 1991).
Let the observed covariates of $a$ and $s$ be $\mathbf{X}^{as}$. Given the distribution of $\epsilon^{as}$, our first-stage estimator recovers the matching parameters $\beta$ in the expression
$U_{as}=\mathbf{X}^{as}\bm{\beta}+\epsilon^{as}.$ (1)
[…]
Second stage (longer-term performance).
In the second stage, post-matching performance, $Y^{as}$, is modeled as
$Y^{as}=\mathbf{X}^{as}\bm{\alpha}+\eta^{as}.$ (2)
Here, $Y$ can be any startup’s performance after it finishes the accelerator program. In particular, the coefficient of gender is our parameter of interest, and it lies in the coefficient vector $\bm{\alpha}$.
Note that the second-stage error term is $\eta^{as}$ correlated with $\epsilon^{as}$. This is because the VC’s decision $Y$ is also correlated with the unobserved match quality $\epsilon^{as}$, which is observed by the accelerator managers and the startups during the admission process. For each potential pair $(a,s)$, we model the random vector $(\epsilon^{as},\eta^{as})$ using a bivariate normal distribution. Without loss of generality in our context, we normalize the variances so that
$\begin{bmatrix}\epsilon^{as}\\ \eta^{as}\end{bmatrix}\sim\mathcal{N}\left(0,\begin{bmatrix}1&\rho\sigma\\ \rho\sigma&\sigma^{2}\end{bmatrix}\right).$ (3)
This specification captures the idea that higher-quality matches—those with larger values of $\epsilon^{as}$—are more likely to produce stronger post-matching outcomes, reflected in higher $\eta^{as}$. By allowing the two error terms to be correlated, the model accounts for the endogeneity that arises when match quality, observed by the matching agents but unobserved by the econometrician, affects both selection and outcomes. The normalization of variances simplifies identification while preserving the structure needed to estimate the correlation parameter $\rho$.
@@ 4.2.1 First Stage: Estimating the Matching Model
[…]
@@
do NOT hold simultaneously. Define the threshold for blocking as
@@
[…]
@@
where $\Phi$ is the c.d.f. of a standard normal distribution due to the assumed marginal distribution of $\epsilon^{as}$ in equation (3). Since this product readily integrates out all unmatched pairs $(a,s^{\prime})$, the overall likelihood of an observed matching $\mu$ is therefore
@@
[…]
@@ 4.2.2 Second Stage: Estimating the Performance of Startup Graduates
In the second-stage analysis, we study ex post outcomes $Y^{as}$ through (2). However, $\eta^{as}$ is not independent of $\bm{X}^{as}$ due to its correlation with $\epsilon^{as}$ as displayed in (3). For example, when $a$ and $s$ are distant from each other, they can still form a match if the realized $\epsilon^{as}$ is large enough. Therefore, among the realized matches, distance correlates with $\epsilon^{as}$. As a result, since $\epsilon^{as}$ correlates with $\eta^{as}$, and $\eta^{as}$ correlates with $\bm{X}^{as}$.
[…]
In other words, suppose that we observed $\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}$, we could use it in (2) to correct for endogeneity problems through a control function approach:
$Y^{as}=\bm{X}^{as}\bm{\alpha}+\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}+\delta$ (4)
where $\delta:=\eta^{as}-\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}$. Since the error term $\delta$ now has expectation zero given $\bm{X}^{as}$ (i.e., $\bm{X}^{as}$ is controlled for in $\bm{X}$), this regression can be estimated with no bias through the ordinary least squares (OLS) method, according to:
Proposition 2.
$\mathbb{E}[\eta^{as}|\mu,\bm{X}]$ is a scalar multiple of $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$, i.e., they are collinear.
Proof.
See Appendix section A.4. ∎
The proposition finds that we can use the observed $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ in place of the unobserved $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ in (4) to obtain an unbiased least squares estimate for $\bm{\alpha}$. We estimate $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ by taking the average of the simulated conditional distribution of $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ from the matching model. The standard errors for $\bm{\alpha}$ is obtained through bootstrapping.
@@ 4.3 Observables
We hand collect several variables relating to startups, accelerators, and their cohorts.
Gender gap.
To examine the gender gap, we define Female Founder as an indicator for whether at least one among a startup’s founders is female. While analyzing all-women founding teams is of interest, the limited number of such cases constrains our analysis. This measure aligns with the divergence trend observed in Pitchbook data, where gender disparities are likely more pronounced for all-women teams (Box and Segerlind, , 2018).
Team member characteristics.
We collect other characteristics of the founding members with information collected from Linkedin and their personal websites. To capture prior entrepreneurship experience, we define No Serial Founder as an indicator equal to one if no founder has prior startup experience. General work experience, proxied by age, serves as a quality signal, which we examine through Average Age of Founding Team. We also include indicators for educational background: At least One Graduate Degree, At least One PhD Degree, and At least One Engr/Sci Degree. Additionally, Founding Team Size captures the number of founders. Industry differences are accounted for using six-digit 2012 NAICS codes.
Startup characteristics.
Longer-established startups differ from newly founded ones, as they are more likely to have a defined business model, customer base, and experienced founders, all of which facilitate investor attraction. To account for this, we include Startup Age, defined as the number of years since founding prior to accelerator admission.
Accelerator characteristics.
We capture accelerator quality variation with log(Cohort Size), while Accelerator Experience measures the years an accelerator has operated before the current cohort. Accelerator in Startup Hubs indicates whether the accelerator is in MA, CA, or NY. [^10] Data limitations prevent controlling for detailed location fixed effects. To account for potential gender-related differences, ‘All-Men’ Accelerator equals 1 if the founders of the accelerator are all male.
Second-stage performance variables.
We measure startup performance using total VC investment received within one year and five years after demo day. Following the literature, we also track survival status (failed or still operating) and acquisition status (acquired or not) by the fifth year (Hockberg et al., , 2007; Gompers et al., , 2010; Ewens and Rhodes-Kropf, , 2015). Since very few accelerator graduates have completed an IPO, we do not consider IPOs as a measure of performance.